OpenStudy (anonymous):

how do i find the slope and y-intercept of the graph 3x – 2y + 8 = 0.

OpenStudy (anonymous):

separate y on LHS you will get what you want

OpenStudy (anonymous):

to get the y intercept substitute the values of x with 0

OpenStudy (anonymous):

to get the slope y=mx+c so you need to transform the equation

OpenStudy (anonymous):

m is the slope

OpenStudy (anonymous):

3x-2y+8=0 2y=3x+8 divide the whole equation to 2 u will get y=3/2(x)+4 so the slope is 3/2

OpenStudy (anonymous):

ok so once i get the y intercept why do i divide it by 2?

OpenStudy (amistre64):

you can go about moving everything around; or, you can just determine the slope and y intercept in 3 moves. 1: y int is the value of y when x=0 2: the slope is the change in y when x changes by 1; so find the value of y when x=1 3: subtract the value when x=0 from the value when x=1 and you have your slope

OpenStudy (anonymous):

as i told you just separate y on LHS you will get y = 3/2x+4 now compare it with y=mx+b to get slope and y-intercept here b= y-intercept and m = slope do it now

OpenStudy (anonymous):

what does LHS mean?

OpenStudy (anonymous):

left hand side

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

now what you got??

OpenStudy (anonymous):

im not sure. and where the 8 go?

OpenStudy (anonymous):

3x – 2y + 8 = 0. subtract 8 from both sides what you get???

OpenStudy (anonymous):

3x-2y=-8?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

now subtract 3x from both sides what you get??

OpenStudy (anonymous):

why substract i thought you you add it? but 2y=-8-3x?

OpenStudy (anonymous):

because 3x on left side has positive sign to cancel it out you have to subtract 3x on both sides

OpenStudy (anonymous):

-2y=-8-3x now divide both sides by -2

OpenStudy (anonymous):

y=4-3/2x

OpenStudy (anonymous):

no its y = 4+3/2x

OpenStudy (anonymous):

oh ok is that the asnwer or is there more to it?

OpenStudy (anonymous):

no you have to do one step now you have a general equation for line in slope intercept form that is y = mx+b here m = slope and b = y intercept can you find the slope and y-intercept now from y = 4+3/2x

OpenStudy (anonymous):

just compare

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now m =?? b= ???

OpenStudy (anonymous):

m=3/2, b=4

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

these are respective values for slope and y-intercept

OpenStudy (anonymous):

ok and that's it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you now i understand it

OpenStudy (anonymous):

you are welcome